OpenStudy (anonymous):

Consider the function f(x)=2x^3−15x^2−36x+8 on the interval [−6,8]. The average or mean slope of the function on this interval is equal to : 38 By the Mean Value Theorem, we know there exists a c in the open interval (−6,8) such that f′(c) is equal to this mean slope. For this problem, there are two values of c that work: The smaller one c1= ? The larger one c2= ? Can someone help me finish this problem?

OpenStudy (lalaly):

find the derivative of f(x) and set it equal to the mean slope and solve for x the x's u get are c1 and c2

OpenStudy (anonymous):

I got 6(x^2-5x-6)=38 then I'm stuck at (x-6)(x+1)=32 I don't know how to solve this...

OpenStudy (lalaly):

no \[6x^2-30x-36=38\]\[6x^2-30x-74=0\]

OpenStudy (lalaly):

now use the quadratic formula

OpenStudy (anonymous):

I get a negative number inside of sqrt thingy...